OpenStudy (anonymous):

help? What is the length (in cm) of the diagonal of a rectangle that has a width of 8 cm and a diagonal of 20 cm?

OpenStudy (anonymous):

L = Length (L-2)= Width By Pythagoras: L^2 + (L -2 )^2 = 10^s L^2 + L^2 - 4L+ 4 = 100 2L^2 - 4L - 96 = 0 Divide by 2 L^2 - 2L - 48 = 0 Simplify: (L+ 6)(L - 8) = 0 Width = 6 Length = 8 (As lengths must be positive)